# Mapping the Spiral of Silence: Surveying Unspoken Opinions in Online Communities

### v1

@@ 3.3 Community Moderation
In this work, we focus on moderation through content removal. We acknowledge that content removal represents only a subset of moderation actions available on Reddit. Subreddit moderators can intervene at the content level, which, in addition to removing content, includes approving content, managing users, changing settings (e.g., rules), or editing flairs / labels (Li, Hecht, and Chancellor 2022). Moderation can occur at the user-level (e.g., deplatforming, shadow-banning) (Rogers 2020; Myers West 2018) or community-level  (Trujillo and Cresci 2022; Chandrasekharan et al. 2022) (e.g., quarantining). We choose to focus on content removal because it is one of the most common moderation interventions and has been shown to influence community members’ participation (Li, Hecht, and Chancellor 2022).
4 Method
@@
Our goal is to measure the spiral of silence across topics and communities. We first need a set of controversial topics that may trigger the spiral of silence. This analysis requires us to identify topics that community members agree are permissible to discuss in their community in principle, but may not feel comfortable sharing in practice. Mining existing posts is insufficient because this data will not capture the opinions that users may be hesitant to share. In other words, starting with existing topics runs the risk of only identifying topics that are so safe as to appear commonly in the community.
@@
Instead, we use large language models (LLMs) to nominate potential controversial topics that are relevant to each community, even if those topics are not posted in the community, then validate those proposed topics with active subreddit members. We then survey community members to capture the rate of self-silencing on those topics.
In this section, we start by discussing the scope of our study and our definition of self-silencing. Then, we provide an overview of our method, describing the topic generation process and survey instrument.
@@
One important distinction we make in this work is between when community members self-silence because they are uncomfortable sharing an opinion versus when an opinion is not considered acceptable to be shared within a community. Our study is constrained only to the former. In other words, we focus only on perspectives that lead to disagreement while still falling within the bounds of the community norms. We do not seek to make any normative claims about what should or should not be shared to a subreddit. Instead, we are interested in what is not being shared even when most community members agree the opinion should be allowed to be discussed.
4.1 Topic Generation
@@
Figure 1: We generate controversial viewpoints specific to a provided community. Using the community name, description, and rules, we prompt an LLM to propose controversial topics. For each topic, we enumerate statements representing different stances on the topic. Finally, the generated topics and viewpoints are validated by community members to ensure they are relevant and fall within the norms of what is accepted in the community.
@@
Our methodological approach involves using an LLM to propose controversial topics and opposing viewpoints for each topic, then surveying actual community members to select viewpoints for our study. While manually selecting topics is the de facto approach when studying the spiral of silence across more than one issue type (Gearhart and Zhang 2018; Lee et al. 2004), manual selection is challenging to perform across many communities on Reddit as each community has its own rules and norms (Fiesler et al. 2018). Deciding what topics are controversial will be imprecise for those without specific knowledge of the community (Hessel and Lee 2019; Chen and Berger 2013). Directly looking at Reddit data is not adequate in this scenario either, as we are interested in the opinions that are unlikely to be shared to the community, rather than those that are already present. In addition to providing a scalable option, LLM-based methods can be used to identify topics that are relevant to the community and might result in self-silencing. However, we cannot assume that all the topics LLMs suggest might be relevant, so we develop a human-plus-AI pipeline that uses an LLM to suggest topics and then has active community members validate the relevancy of these proposals.
@@
Topic generation is a three-step process (see Fig. 1). We provide a community name, description, and rules as input to an LLM (GPT-4) to propose 20 potentially controversial topics for the provided community. “Controversial” is defined as being likely to cause disagreement between community members. Then, for each proposed topic, the LLM generates viewpoints representing different sides of the issue using few-shot prompting techniques. The viewpoints must follow the rules of the subreddit, which are again provided to the model. Finally, the viewpoints are manually validated by active community members.
Proposing Controversial Political Topics and Viewpoints
@@
First, we must select which subreddits to provide as input to the LLM. We focus on subreddits that focus on political and social issues, since Noelle-Neumann (1974) posits that morally laden subjects typically trigger the spiral of silence. To identify these communities, we start with the 2,040 subreddits on r/ListOfSubreddits and sample the 50 hottest submissions in each. We then filter for the following criteria: (1) the majority of submissions are in English as classified by langdetect (Shuyo 2010); (2) $\geq 80\%$ of submissions are classified to contain political content; and (3) the subreddit has over 100K members. Filtering leaves us with 23 political subreddits.
Since we seek to compare variation in opinion expression on topics based on community-level differences, we want to identify a single, static set of topics that are relevant across all 23 subreddits. We select r/politics and r/worldnews as the communities to input to the LLM based on their size and high proportion of political content. Using our pipeline, we generate 40 topics (i.e., 20 per subreddit) with opposing viewpoints for each topic.
Validating Generated Topics and Viewpoints
Next, we validate the generated topics and viewpoints. Starting with the 40 topics from the prior step, we check the generated outputs for comprehension. Four had incomplete viewpoints (i.e., GPT-4 only provided viewpoints representing one side) and three were either worded in a confusing manner or presented viewpoints that did not present opposite stances, leaving us with 33 topics.
We further validated the remaining topics by working with Prolific participants who self-reported as active members of r/politics or r/worldnews. First, participants completed a screening quiz to confirm their knowledge about the subreddits. In the quiz, participants selected which post titles were likely to appear in their selected subreddit from a set of three titles—two from their selected subreddit and one from another political subreddit (see Appendix A.3 for examples). If they identified all the correct post titles, participants qualified for the survey. Participants were asked whether a given topic is relevant to the subreddit.

### v2

@@ 3.3. Community Moderation
In this work, we focus on moderation through content removal. We acknowledge that content removal represents only a subset of moderation actions available on Reddit, but choose to focus on content removal because it is one of the most common moderation interventions and has been shown to influence community members’ participation (Li et al., 2022).
4. Method
@@
Our goal is to measure the spiral of silence across topics and communities. We first need a set of controversial topics that may trigger the spiral of silence. This analysis requires us to identify topics that community members agree are permissible to discuss in their community in principle, but may not feel comfortable sharing in practice. Crawling existing posts from these communities, unfortunately, is insufficient: these posts will not reflect the opinions that users may be hesitant to share in the first place. In other words, starting with existing topics runs the risk of only identifying topics that are so safe as to appear commonly in the community. Alternatively, while asking participants to list instances when they have self-silenced is possible, this elicitation process may be prone to recall or desirability bias and yield idiosyncratic or off-scope topics falling outside the theory’s bounds.
@@
Thus, we use a hybrid human-AI method that draws on large language models (LLMs) to nominate potential controversial topics that are appropriate to each community, even if those topics are not posted in the community, and validate those proposed topics with active subreddit members. We then survey community members to capture the rate of self-silencing on those topics. In this section, we start by discussing the scope of our study and our definition of self-silencing. Then, we provide an overview of our method, describing the topic generation process and survey instrument. This study was approved by Stanford University’s Institutional Review Board.
@@
One important distinction we make in this work is between when community members self-silence because they are uncomfortable sharing an opinion versus when an opinion is not considered acceptable to be shared within a community. Our study is constrained only to the former. In other words, we focus only on perspectives that lead to disagreement while still falling within the bounds of the community norms. We do not seek to make any normative claims about what should or should not be shared to a subreddit. Instead, we are interested in what is not being shared, even when most community members agree the opinion should be allowed to be discussed.
4.1. Topic Generation
@@
Figure 2. We generate controversial viewpoints specific to a provided community. Using the community name, description, and rules, we prompt an LLM to propose controversial topics. For each topic, we enumerate statements representing different stances on the topic. Finally, the generated topics and viewpoints are validated by community members to ensure they are relevant and fall within the norms of what is accepted in the community.
Overview of the topic generation pipeline. Overview of the topic generation pipeline. The diagram illustrates how subreddit information (name, description, and rules) is provided to an LLM, which generates potentially controversial topics and opposing viewpoints. In the final step, community members then validate these topics to ensure relevance and appropriateness.
@@
Our methodological approach involves using an LLM to propose controversial topics and opposing viewpoints for each topic, then surveying actual community members to select viewpoints for our study, which provides several advantages for studying the spiral of silence in our setting. First, while manually selecting topics is the de facto approach when studying the spiral of silence across more than one issue type (Gearhart and Zhang, 2018; Lee et al., 2004), manual selection may introduce experimenter bias and may be impractical across a large number of diverse yet specific communities on Reddit, each with its own rules and norms (Fiesler et al., 2018). Deciding what topics are controversial will be imprecise for those without specific knowledge of the community (Hessel and Lee, 2019; Chen and Berger, 2013). Second, directly looking at Reddit data is not adequate in this scenario either, as we are interested in the opinions that may be less likely to be shared to the community, rather than those that are already present. Thus, using mined data from online only gives us a window into the viewpoints that community members feel comfortable sharing. By contrast, using an LLM allows us to nominate potential topics that are relevant to the community and might result in self-silencing, including perspectives that may be underrepresented or missing in the current discourse online. Prior work has also shown that LLM-based topic generation produces more coherent and interpretable topics than traditional approaches such as LDA (Pham et al., 2024; Kaur and Wallace, 2024). We then ask active community members to validate the relevancy and appropriateness of these proposals, selecting topics from this filtered set to include in our study.
4.1.1. Proposing Controversial Topics and Viewpoints
@@
Topic generation is a three-step process (see Fig. 2). We provide a community name, description, and rules as input to an LLM (GPT-4) to propose 20 potentially controversial topics for the provided community. We provide in our prompt that we define “controversial” as being likely to cause disagreement between community members. Then, for each proposed topic, the LLM generates two viewpoints representing different sides of the issue using few-shot prompting techniques (see Appendix A.2 for prompts). The viewpoints must follow the rules of the subreddit, which are again provided to the model. Finally, as we discuss in detail in Sec. 4.2, viewpoints are manually validated to ensure that they are comprehensible and relevant to the respective subreddit.
4.1.2. Selecting Political Subreddits
@@
To run our topic generation pipeline, we must select which subreddits to provide as input to the LLM. For this work, we focus on subreddits that focus on political and social issues, since Noelle-Neumann (1974) posits that morally laden subjects typically trigger the spiral of silence. To identify these communities, we start with the 2,040 subreddits on r/ListOfSubreddits and sample the 50 hottest submissions in each as of January 2024 accessed via the Reddit API. We then filter for the following criteria: (1) the majority of submissions are in English as classified by langdetect (Shuyo, 2010); (2) $\geq 80\%$ of submissions are classified to contain political content using an LLM classifier on the post title following an approach from prior work (Piccardi et al., 2025) (see Appendix A.3 for prompt); and (3) the subreddit has over 100K members. Filtering leaves us with 23 political subreddits.